OpenStudy (yoyo2443):

Solve the linear system using substitution:

OpenStudy (yoyo2443):

\[-x+\frac{ 2 }{ 9 }y=1\] \[9x+3y=6\] I've been trying to solve this for a while now, but can't seem to get the answer, please help!

OpenStudy (ajspeller):

Multiply each term in the first equation by 9. Then add both of the equations together. This will eliminate the x variable. You can then solve for y. OR Solve one of the equations for y. Then substitute the expression for y that you just found into the other equation. OR Solve one of the equations for x. Then substitute the expression for x that you just found into the other equation. AND View this playlist for examples on how to solve those types of problems. https://www.youtube.com/playlist?list=PLsX0tNIJwRTwHqLbfCQTZdVUmqz1KBHgv I hope this helps.

563blackghost (563blackghost):

To solve lets change the second equation into slope-intercept form.... First we would subtract by 9x to get it to the other side... \(\LARGE{(9x-9x)+3y=-9x+6 \rightarrow 3x=-9x+6}\) Next we would divide by 3 to isolate y.... \(\huge{\frac{3y}{3}=\frac{-9x}{3}+\frac{6}{3}}\) What would that equal?

OpenStudy (yoyo2443):

\[y=-3+2\]

563blackghost (563blackghost):

Your missing the x where the 3 is so it would be...but Correct :) \(\huge{y=-3x+2}\) Now we know the equation for y...we can now input y into the first equation... \(\huge{-x+\frac{2}{9}(-3x+2)=1}\) In which we would simplify.

OpenStudy (yoyo2443):

oh right, I forgot the x my bad

563blackghost (563blackghost):

All's good just make sure to remember next time ;)

563blackghost (563blackghost):

So now we solve the equation above...we would first distribute \(\Large{\frac{2}{9}}\) into the equation in parenthesis.... \(\huge{-x+((-3 \times \frac{2}{9})x+(2 \times \frac{2}{9}))=1}\) What would that equal?

OpenStudy (yoyo2443):

\[-x-\frac{ 6 }{ 9 }x+\frac{ 4 }{ 9 }=1\]

563blackghost (563blackghost):

Close but not quite. You need to simplify the fraction... \(\huge{-x-\frac{2}{3}x+\frac{4}{9}=1}\) Next combine like terms....

563blackghost (563blackghost):

What would.. \(\huge{-x-\frac{2}{3}x=?}\)

OpenStudy (yoyo2443):

\[-\frac{ 5 }{ 3 }x\] ?

563blackghost (563blackghost):

Correct :) \(\huge{-\frac{5}{3}x+\frac{4}{9}=1}\) Now we need to get the real numbers to one side so we would subtract \(\Large{\frac{4}{9}}\) on both sides... \(\huge{-\frac{5}{3}x(\frac{4}{9}-\frac{4}{9})=1-\frac{4}{9}}\) What does that simplify too?

OpenStudy (yoyo2443):

\[-\frac{ 5 }{ 9 }x=\frac{ 5 }{ 9 }\]

563blackghost (563blackghost):

Correct :) except the fraction with the variable will stay as `-5/3x`.... Now we would isolate the variable by dividing `-5/3`.... \(\huge{(-\frac{5}{3}\div \frac{5}{3})x=\frac{5}{9} \div \frac{5}{3}}\)

OpenStudy (yoyo2443):

oh ugh im sorry I forgot that too

563blackghost (563blackghost):

What do you mean?

OpenStudy (yoyo2443):

I forgot that it was 5/3 and not 5/9

563blackghost (563blackghost):

Ooo yea...all's good though :)

OpenStudy (yoyo2443):

anyways \[x=\frac{ 1 }{ 3 }\] ?

563blackghost (563blackghost):

Correct :) Though I forgot to put a negative sign by `5/3` so that was my mistake so.... \(\LARGE{x=-\frac{1}{3}}\) This is the x-value of the answer now we need to find the y-value... Now we would input x into the second equation.... \(\huge{y=-3(-\frac{1}{3})+2}\) Simplify ;)

563blackghost (563blackghost):

Were you able to find what y equals?

OpenStudy (yoyo2443):

\[y=3\]

563blackghost (563blackghost):

Correct :) So the solution of the linear system is....\(\LARGE{(-\frac{1}{3},3)}\) Nicely done!

OpenStudy (yoyo2443):

Ah okay, thank you!

OpenStudy (yoyo2443):

I was really breaking my head over this, thanks a lot :)

563blackghost (563blackghost):

Your very welcome :) If you ever need help in the future and see me on just tag me and Ill come as soon as I can (thats if im not busy) And I dont know if anyone sent you a welcome message when you joined soo.... \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) \(\Large{By:thomaster}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct

OpenStudy (yoyo2443):

Oh no one had haha, thank you!

563blackghost (563blackghost):

:)